OpenStudy (anonymous):

what is an equation for a line that has a slope of 2/3 passing through (5,-7) ?

OpenStudy (geometry_hater):

Input the x and y and your slope into an equatio -7 = 2/3 (5) + B Solve for b

OpenStudy (geometry_hater):

Y= MX+ B

OpenStudy (anonymous):

or in point-slope form: \(\Large y-(-7)=-\frac{2}{3}(x-5) \) simplify....